9.1 Let x be a random variable representing dividend yield of bank stocks. We may assume...

9.1

Let x be a random variable representing dividend yield of bank stocks. We may assume that x has a normal distribution with σ = 2.7%. A random sample of 10 bank stocks gave the following yields (in percents). 5.7 4.8 6.0 4.9 4.0 3.4 6.5 7.1 5.3 6.1

The sample mean is x = 5.38%. Suppose that for the entire stock market, the mean dividend yield is μ = 4.1%. Do these data indicate that the dividend yield of all bank stocks is higher than 4.1%? Use α = 0.01.

(a) What is the level of significance?

What is the value of the sample test statistic? (Round your answer to two decimal places.)

(c) Find (or estimate) the P-value. (Round your answer to four decimal places.)
Answer #1

a)

0.01 is the level of significanc e

H0 : mu = 4,1
Ha: mu > 4.1

This is a one tailed test

test statistics;

z = (xbar -mu)/(sigma/sqrt(n))
= ( 5.38 - 4.1)/(2.7/sqrt(10))
= 1.50

b)

p value = 1- P(z< 1.50)
= 1- 0.9331
= 0.0669

p value = 0.0669

c)

Fail to reject H0 as p value > 0.01 significanc elevel